ANLA Citation List

HoOtto92 [1] S. Holmgren and K. Otto, Iterative solution methods and preconditioners for block-tridiagonal systems of equations, SIAM J. Matrix Anal. Appl., 13 (1992), pp. 863–886.

Larsson99 [2a] E. Larsson, A domain decomposition method for the Helmholtz equation in a multilayer domain, SIAM J. Sci. Comput., 20 (1999), pp. 1713–1731.

OttoLa99 [2b] K. Otto and E. Larsson, Iterative solution of the Helmholtz equation by a second-order method, SIAM J. Matrix Anal. Appl., 21 (1999), pp. 209–229.

HoOtto94 [2c] S. Holmgren and K. Otto, Semicirculant preconditioners for firstorder partial differential equations, SIAM J. Sci. Comput., 15 (1994), pp. 385–407.

HeOtto96 [2d] L. Hemmingsson and K. Otto, Analysis of semi-Toeplitz preconditioners for first-order PDEs, SIAM J. Sci. Comput., 17 (1996), pp. 47– 64.

AxNey03a [2e] O. Axelsson and M. Neytcheva, Preconditioning methods for linear systems arising in constrained optimization problems, Numer. Linear Algebra Appl., 10 (2003), pp. 3–31.